# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

### v1

@@ II-B Related Work: Perplexity for Code Authorship Attribution
Detecting LLMgCode from HaCode can be regarded as a particular code authorship attribution problem [51, 52, 53, 54] in which the authors are either LLMs or humans. Code authorship attribution fills in the missing or inaccurate author information of source code fragments, which is beneficial for software maintenance [55, 56, 57] and software quality analysis [58, 59, 60, 61]. For example, code authorship attribution helps detect code plagiarism [62, 63] and trace malware provenance [64]. Moreover, detecting LLMgCode from HaCode increases the authenticity and integrity of software [14], reducing risks such as security vulnerabilities and copyright infringement in the software supply chain [15].
@@
Perplexity is originally one of the most popular methods for detecting AI-generated content (including but not limited to LLMgCode) [23, 24, 25]. For example, Mitchell et al. [24] proposed DetectGPT to detect LLM-generated text. DetectGPT utilizes the log probabilities to calculate the perplexity. The authors evaluated DetectGPT using the AUC metric on 500 LLM-generated samples and 500 human-written samples. In addition, they evaluated the generalization capability of DetectGPT by alternately using GPT-J-6B, GPT-Neo-2.7B, and GPT-2-1.5B as the code generation model and the perplexity calculation model. Su et al. [25] proposed two methods to calculate the perplexity. The first is Log-Likelihood Log-Rank Ratio (LRR) and the second is Normalized Perturbed Log-Rank (NPR). The authors evaluated LRR and NPR using the AUC metric on 300 LLM-generated and 300 human-written samples. They also compared LRR and NPR against baselines in terms of inference efficiency.
@@
With the increasing popularity of LLMgCode, Perplexity has been commonly used for detecting LLMgCode from HaCode [14, 26]. For example, Xu and Sheng [26] proposed a method using the variant of probability to calculate the perplexity. They evaluated their methods using the AUC, FPR, and TPR metrics on 5,214 LLM-generated and 5,214 human-authored code snippets. They also evaluated the influence of models used to calculate perplexity. Shi et al. [14] proposed a method namely DetectCodeGPT to detect LLMgCode from HaCode. DetectCodeGPT calculates the perplexity based on NPR. They evaluated DetectCodeGPT using the AUC metric on 500 LLM-generated and 500 human-authored code snippets. They also evaluated the generalization capability of DetectCodeGPT by alternately using Incoder-1B, Phi-1B, StarCoder-3B, WiardCoder-3B, CodeGen2-3.7B, and GodeLlama-7B as the code generation model and the perplexity calculation model.
@@
Overall, Perplexity is a popular method for detecting LLMgCode from HaCode but its efficacy evaluation is limited, which does not reflect its effects well in practice. In this article, we carry out a large-scale investigation on Perplexity from multiple realistic aspects.
III Motivating Examples
@@
III-A Code Authenticity
Although Perplexity has been demonstrated useful for detecting LLMgCode from HaCode, most of the LLMgCode is generated by LLMs with relatively a few parameters, such as CodeLlama-7B [14].
Figure 2: An example of the code generated by CodeLlama-7B and the code written by human
@@ III-A Code Authenticity
Figure 2 presents examples of LLMgCode and HaCode. The prompt indicates that the function to be generated is to make a git commit when the template files are modified. The prompt is the input of CodeLlama-7B (shown in the first block). The LLMgCode is the solution generated by CodeLlama-7B (shown in the second block), and the HaCode is the solution provided by a human developer (shown in the third block). The LLMgCode is not an acceptable solution and differs significantly from the HaCode. Specifically, the LLMgCode contains only repeated comments. Although it is easy to detect such LLMgCode from HaCode, few developers use such a code generation model like CodeLlama-7B as it is far from addressing real-world problems.
@@
When Perplexity uses the same model for code generation and detection, it might achieve low perplexities of LLMgCode. This is because the model would assign a high prediction probability to each token when it perceives the code generated by itself. However, in real-world scenarios, the detection model does not have clues to know which model is used to generate code early. Moreover, these LLMs used for generating code have hundreds of billions of parameters, making them challenging to deploy locally to calculate perplexities.
III-B Language Diversity
@@
Perplexity is evaluated in limited programming languages so far. However, the code stylometry of different programming languages varies greatly. Figure 3 presents two pieces of code generated by GPT-3.5, both of which are to calculate the median of all subarrays of a given array, and then calculate the median of these medians.
Figure 3: An example of code achieving the same task in Python and C
@@
As illustrated in Figure 3, the code stylometries differ in Python and C, particularly in their logic and scale. Python has a rich standard library and built-in functions. For example, “statistics.comedian” is a tool library in Python. It can directly calculate the median without the need for manual sorting and selection. while C does not have a direct median calculation function, so we need to use “qsort” to sort and manually calculate the median. In fact, many high-level functions must be manually implemented in C, such as inserting or deleting elements in arrays. However, such functions can be automatically handled by the built-in methods of lists in Python. In addition, Python is a dynamically typed programming language that does not require declaring variable types, making the code more concise. While C is a statically typed programming language where each variable needs to declare its type. Moreover, manipulating elements such as list slicing is easy in Python. For example, “a[i: j+1]” returns subarrays without manual calculation and traversal. While C requires manually managing memory and arrays, which means that developers need to handle each element carefully.
@@
Due to differences such as the built-in library support and design principles, code written in different programming languages exhibits variations. In addition to programming languages, the code used to solve puzzles of different difficulty and the code of different scales exhibits variations in terms of logic and complexity.
III-C Generalization Capability
@@
There are many LLMs available for generating code, such as GPT-4o, Copilot, and Gemini. Due to differences in training data, model architecture, optimization objectives, generation strategies, etc., the code generated by these LLMs is different. Figure 4 presents two Java code snippets generated by GPT-4o and Gemini solving the same programming puzzle. The puzzle is a dynamic programming problem that is similar to the unbounded knapsack problem. This problem can be described as follows. Given the monster’s initial health $H$ and $N$ different spells, each spell has a fixed damage $A_{i}$ and a magic point cost $B_{i}$. The objective is to reduce the monster’s health to 0 or below by casting spells while minimizing the total magic points consumed.
@@
Figure 4: An example of code achieving the same task generated by GPT-4o and Gemini
@@
As shown in Figure 4, although both pieces of code aim to reduce the monster’s health to 0 with minimal loss, the two LLMs have different implementations. GPT-4o starts to generate code from the perspective of damage, aiming to achieve total damage $H$ with the minimum cost. Gemini starts to generate code from the perspective of the monster’s health, aiming to reduce the monster’s health to 0 with the minimum cost. The former is to intuitively accumulate damage in the forward direction, and the latter is to reduce the monster’s health in the reverse direction.
@@
In addition to the above differences, the abstraction levels of implementations are different. The solution generated by Gemini is detailed and complex. Gemini uses a two-dimensional array ‘dp’ to simultaneously record health values and spell combinations. GPT-4o only uses a one-bit array ‘dp’ to record the minimum cost of achieving each damage value. When updating the state, Gemini considers all spell combinations in each update, making it more complex; while GPT-4o directly updates the state by traversing damage values and spell methods. In addition, Gemini considers boundary conditions and special cases, such as returning an outlier of “-1” when the target cannot be reached. Unfortunately, GPT-4o fails to handle these boundary conditions, reducing its robustness.
@@
In short, GPT-4o focuses on the core logic; while Gemini focuses on practical usages, boundary conditions, and robustness on this example. This example motivates us to investigate the generalization capability of Perplexity when the code is generated by different LLMs. In addition to generalization, the efficiency of Perplexity is important, as it influences the usage scenarios and user experience of Perplexity.
IV Research Methodology
This section elaborates on the research methodology of detecting LLMgCode from HaCode. shows an overview of our research design and process.
IV-A Research Questions
@@
Existing work focuses on evaluating the detection accuracy (effectiveness) of Perplexity, yet there are many aspects should be considered when using it in practice. Toward realistic detection, we intend to answer three research questions (RQs) through this investigation.
@@
    Is Perplexity an effective method for detecting LLMgCode?: Language ($\text{{RQ}}_{\textbf{1.1}}$), Difficulty ($\text{{RQ}}_{\textbf{1.2}}$), and Scale ($\text{{RQ}}_{\textbf{1.3}}$)
@@
    Detection accuracy is crucial for detecting LLMgCode, while it might not be stable between situations, e.g., an effective method for C++ may not suit Python. Therefore, we propose $\text{RQ}_{1}$ to compare Perplexity against state-of-the-art methods regarding effectiveness from three sub-aspects: (1) programming languageof the implementation: C, C#, C++, Go, Java, Python, Ruby, and Rust; (2) degree of difficultyof the puzzle: easy, medium, and hard; and (3) scale of solutionto the puzzle: small, middle, and large.
@@
    Developers utilize various LLMs to assist in programming. Due to differences in training data, model architecture, and generation strategies, the code generated by these LLMs can vary significantly. An optimal detection method should be capable of detecting the code generated by different LLMs. In $\text{RQ}_{2}$, we investigate the generalization capability of Perplexity in detecting LLMgCode.
  • $\text{{RQ}}_{\textbf{3}}$:
    Is Perplexity an efficient method for detecting LLMgCode?
    Motivation:
    Detection speed (efficiency) influences the user experience. According to the detection mechanism, Perplexity applies perturbation strategies to generate dozens of perturbation samples and uses an LLM to calculate the probability or rank of each token for both the original sample and perturbation samples [14, 25]. Such a two-phase detection process might be time-consuming. In $\text{RQ}_{3}$, we compare Perplexity with baselines from the perspective of inference efficiency.
IV-B Data Preparation
Although there are several dataset [14, 15, 19] for detecting LLMgCode from HaCode, none of them is suitable to address all research questions. Therefore, we developed a new one based on CodeNet [27]. CodeNet was originally proposed by IBM and built to promote AI technologies in the domain of code. Since then, CodeNet has been used for many studies [26, 65, 66, 67, 68, 69].
Figure 5: An example of a programming puzzle in CodeNet
@@ IV-B Data Preparation
As shown in Figure 5, a programming puzzle provided by CodeNet has seven basic elements: (1) statement, (2) input, (3) output, (4) constraint, (5) difficulty score, (6) sample input, and (7) sample output. The statement describes what the puzzle is. The input and output describe the requirements for input and output, respectively. The constraint describes the specific constraints on the variables. The difficulty score measures the difficulty level of the puzzle. Finally, the sample input and sample output provide the input and output of a specific example, respectively. Based on CodeNet, we developed our dataset through the following five steps.
IV-B1 Step-1: Problem Selection
To build a standard dataset for evaluation, we excluded the programming problem whose (1) statement is not in English as it is the most popular language; (2) statement contains images or tables as they can hardly be extracted and input into LLMs; (3) solutions are all less than 30 LOC (Lines Of Code) as they are mainly for trivial puzzles; (4) solutions are incomplete regarding all the eight programming languages as they constitute the HaCode to be evaluated in $\text{RQ}_{1.1}$; (5) structure is not complete, i.e., statement, input, output, constraints, difficulty, and samples ($\geq 2$). After employing these criteria, our dataset includes 729 programming puzzles.
IV-B2 Step-2: HaCode Collection
We randomly collected two solutions that were provided by human developers from CodeNet for each puzzle regarding each of the eight programming languages.
IV-B3 Step-3: LLMgCode Generation
We generated LLMgCode for programming puzzles using OpenAI GPT-3.5 (gpt-3.5-turbo-0125). The prompts for code generation are shown in Figure 6. There are two solutions for each puzzle regarding each of the eight programming languages. Moreover, to address $\text{RQ}_{2}$, we used the prompts from Figure 6 and input them into GPT-4o, Gemini, and Llama, with each model generating 500 solutions. Note that the following two steps are performed on the code solutions that have been generated by GPT-3.5.
@@ IV-B3 Step-3: LLMgCode Generation
Figure 6: The prompts used for generating LLMgCode solutions for programming puzzles
IV-B4 Step-4: Classification According to the Difficulty of Puzzles
We classified the puzzles and solutions into three groups based on difficulty scores of puzzles: easy for $\leq 200$, medium for $>200$ and $\leq 500$, and hard for $>500$.
IV-B5 Step-5: Classification According to the Scale of Solutions
We classified the puzzles and solutions into three groups based on the scale of solutions: small for $\leq 20$, medium for $>20$ and $\leq 50$, and large for $>50$.
IV-C Studied Methods
We compared Perplexity against feature-based and pre-training-based methods aimed at detecting LLMgCode from HaCode as the experimental subjects.
IV-C1 Feature-based Methods
Two feature-based methods are as follows.
  • Idialu et al. [19] that leverages a combination of layout, syntactic, and lexical features in Python source code as the input. This method uses XGBoost [70] to construct the detection method.

### v2

@@ 2.2. Related Work: Perplexity for Code Authorship Attribution
Detecting LLMgCode from HaCode can be regarded as a code authorship attribution problem (Abuhamad et al., 2018; Dauber et al., 2018; Bogomolov et al., 2021) in which the authors are either LLMs or humans. Code authorship attribution fills in the missing or inaccurate author information of source code fragments, which is beneficial for software maintenance (Girba et al., 2005; Anvik et al., 2006; Fritz et al., 2010) and software quality analysis (Rahman and Devanbu, 2011; Bird et al., 2011; Yin et al., 2011; Thongtanunam et al., 2016). For example, code authorship attribution helps detect code plagiarism (Burrows and Tahaghoghi, 2007; Kothari et al., 2007) and trace malware provenance (Caliskan-Islam et al., 2015). Moreover, detecting LLMgCode from HaCode increases the authenticity and integrity of software (Shi et al., 2025), reducing a variety of risks such as security vulnerabilities and copyright infringement in the software supply chain (Bukhari et al., 2023). In addition, LLMgCode is typically not considered original (Nguyen et al., 2024), thus detecting LLMgCode is helpful in scenarios that require individual work, such as programming contests (Idialu et al., 2024; Nguyen et al., 2024). This scenario also asks for a high detection efficiency, as examination systems need to provide quick feedback after submission.
@@
Perplexity is originally one of the major methods used for detecting AI-generated content (including but not limited to LLMgCode) (Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023). For example, Mitchell et al. (2023) proposed DetectGPT to detect LLM-generated text. Technically, DetectGPT utilizes the log probabilities to calculate the perplexity. The authors evaluated DetectGPT using the AUC metric on 500 LLM-generated samples and 500 human-written samples. In addition, they evaluated the generalization capability of DetectGPT by alternately using GPT-J-6B, GPT-Neo-2.7B, and GPT-2-1.5B as the code generation model and the perplexity calculation model. Su et al. (2023) proposed two methods to calculate the perplexity. The first is Log-Likelihood Log-Rank Ratio (LRR) and the second is Normalized Perturbed Log-Rank (NPR). The authors evaluated LRR and NPR using the AUC metric on 300 LLM-generated and 300 human-written samples. They also compared LRR and NPR against baselines in terms of detection speed.
@@
With the increasing popularity of LLMgCode, Perplexity has been commonly used in detecting LLMgCode from HaCode (Shi et al., 2025; Xu and Sheng, 2024). Xu and Sheng (2024) proposed a method using a variant of probability to calculate the perplexity. They evaluated their methods using the AUC, FPR, and TPR metrics on 5,214 LLM-generated and 5,214 human-authored code snippets. They also evaluated the influence of models used to calculate perplexity. Shi et al. (2025) proposed a method namely DetectCodeGPT to detect LLMgCode from HaCode. DetectCodeGPT calculates the perplexity based on NPR. They evaluated DetectCodeGPT using the AUC metric on 500 LLM-generated and 500 human-authored code snippets. They also evaluated the generalization capability of DetectCodeGPT by alternately using Incoder-1B, Phi-1B, StarCoder-3B, WiardCoder-3B, CodeGen2-3.7B, and GodeLlama-7B as the code generation model and the perplexity calculation model. These studies focus on proposing new perplexity-based detection methods and evaluating overall detection accuracy. As a result, they only partially consider other detection methods and pay limited attention to broader evaluation aspects, leading to limitations in both evaluation objectives and dimensions. To address this gap, this study conducts a systematic evaluation of feature-based, pre-training-based, and perplexity-based methods from multiple perspectives, including detection accuracy (language, difficulty, scale), detection speed, and generalization capability.
@@
Overall, Perplexity is a popular method for detecting LLMgCode from HaCode but its evaluation is limited. In this article, we conduct a large-scale investigation on the efficacy of Perplexity from multiple aspects.
3. Motivating Examples
@@
3.1. Code Authenticity
Although Perplexity has been demonstrated useful in detecting LLMgCode, the LLMgCode used in previous work is generated by “tiny” LLMs, such as CodeGen2-3.7B (Shi et al., 2025). Their effectiveness for larger code generators is unknown.
Figure 2. An example of the LLMgCode (CodeGen2-3.7B) and HaCode
@@ 3.1. Code Authenticity
Figure 2 presents examples of LLMgCode and HaCode. The prompt indicates that the function to be generated is to synchronize the file state across editor stacks to prevent unnecessary reloads after a save. The prompt is the input of CodeGen2-3.7B (shown in the first block). The LLMgCode is the solution generated by CodeGen2-3.7B (shown in the second block), and the HaCode is the solution provided by a human developer (shown in the third block). The LLMgCode is an incorrect implementation of this functionality and differs significantly from HaCode. In fact, the LLMgCode only returns when the file was not saved under a new name, which is inconsistent with the intended functionality. While HaCode avoids unnecessary reloads by updating the file state: it locates the original filename and synchronizes its metadata, including the creation status, filename, and newest modified time. Although it is easy to detect such LLMgCode from HaCode, few developers use such LLMgCode for their code solutions, as it is far from addressing real-world problems.
@@
When Perplexity uses the same model for code generation and detection, it achieves low perplexities of LLMgCode. This is because the model would assign a high prediction probability to each token when it perceives the code generated by itself. However, in real-world scenarios, the detection model does not have clues to know which model is used to generate code early. Moreover, these LLMs used for generating code have hundreds of billions of parameters, making them challenging to deploy locally to calculate perplexities.
3.2. Language Diversity
@@
Perplexity has been evaluated in limited programming languages so far. However, the code stylometry of different programming languages varies greatly. Figure 3 presents two pieces of code generated by GPT-3.5, aiming to calculate the median of all subarrays of a given array and then calculate the median of these medians.
Figure 3. An example of code achieving the same task in Python and C
@@
As illustrated in Figure 3, the code stylometries differ in Python and C, particularly in their logic and scale. Python has rich built-in libraries. For example, “statistics.comedian” is a tool library in Python. It can directly calculate the median without the need for manual sorting and selection. While C does not have a direct median calculation function, therefore, we use “qsort” to sort and manually calculate the median. In fact, many high-level functions must be manually implemented in C, such as inserting or deleting elements in arrays. However, such functions can be automatically handled by the built-in methods of lists in Python. In addition, Python is a dynamically typed programming language that does not require declaring variable types, making the code more concise. While C is a statically typed programming language where each variable type needs to be declared. Moreover, manipulating elements such as list slicing is easy in Python. For example, “a[i: j+1]” returns subarrays without manual calculation and traversal. While C requires manually managing memory and arrays, which means that developers need to handle each element carefully.
@@
Due to differences such as the built-in library support and design principles, code written in different programming languages is distinguishable. In addition to programming languages, the code used to solve puzzles of different difficulty and the code of different scales are distinguishable in terms of logic and complexity.
3.3. Generalization Capability
@@
There are many LLMs available for generating code, such as GPT, Copilot, and Gemini. Due to differences in training data, model architecture, optimization objectives, generation strategies, etc., the code generated by these LLMs is different. Figure 4 presents two Java code snippets generated by GPT-4o and Gemini-1.0 solving the same programming puzzle. The puzzle is a dynamic programming problem that is similar to the unbounded knapsack problem. This problem can be described as follows. Given the monster’s initial health $H$ and $N$ different spells, each spell has a fixed damage $A_{i}$ and a magic point cost $B_{i}$. The objective is to reduce the monster’s health to 0 or below by casting spells while minimizing the total magic points consumed.
@@
Figure 4. An example of code achieving the same task generated by GPT-4o and Gemini-1.0
@@
As shown in Figure 4, although both pieces of code aim to reduce the monster’s health to 0 with minimal loss, the two LLMs have different implementations. GPT-4o starts to generate code from the point of view of damage, with the aim of achieving total damage $H$ at the minimum cost. Gemini-1.0 starts to generate code from the perspective of the monster’s health, aiming to reduce the monster’s health to 0 with the minimum cost. The former is to intuitively accumulate damage in the forward direction, and the latter is to reduce the monster’s health in the reverse direction.
@@
In addition to the differences mentioned above, the abstraction levels are different. The solution generated by Gemini-1.0 is detailed and complex. Gemini-1.0 uses a two-dimensional array ‘dp’ to simultaneously record health values and spell combinations. GPT-4o uses only a one-bit array ‘dp’ to record the minimum cost of achieving each damage value. When updating the state, Gemini-1.0 considers all spell combinations in each update, making it more complex, while GPT-4o directly updates the state by traversing damage values and spell methods. In addition, Gemini-1.0 considers boundary conditions and special cases, such as returning an outlier of “-1” when the target cannot be reached. Unfortunately, GPT-4o fails to handle these boundary conditions, reducing its robustness.
@@
In short, GPT-4o focuses on core logic, while Gemini-1.0 focuses on practical usages, boundary conditions, and robustness regarding this example. This example motivates us to investigate the generalization capability of Perplexity when the code is generated by different LLMs. In addition to generalization, the detection speed of Perplexity is important, as it influences user experience.
4. Research Methodology
This section elaborates on the experimental designs to investigate the efficacy of Perplexity in detecting LLMgCode.
4.1. Research Questions
@@
Existing work focuses on evaluating the detection accuracy of Perplexity, however, there are many aspects to consider when using it in practice. In this article, we investigate the efficacy of Perplexity in detecting LLMgCode in a wider range of realistic evaluation settings.
@@
    Is Perplexity an effective method in detecting LLMgCode?: Language ($\text{{RQ}}_{\textbf{1.1}}$), Difficulty ($\text{{RQ}}_{\textbf{1.2}}$), and Scale ($\text{{RQ}}_{\textbf{1.3}}$)
@@
    Detection accuracy is crucial in detecting LLMgCode, while it might not be consistently good between situations, e.g., an effective method for C++ might not be suitable for Python. We propose $\text{RQ}_{1}$ to compare Perplexity against state-of-the-art methods regarding detection accuracy from three sub-aspects: (1) programming languageof the implementation: C, C#, C++, Go, Java, Python, Ruby, and Rust; (2) degree of difficultyof the puzzle: easy, medium, and hard; and (3) scale of solutionto the puzzle: small, middle, and large.
@@
    Is Perplexity an efficient method in detecting LLMgCode?
    Motivation:
    Detection speed affects user experience in just-in-time scenarios such as time-limited programming contests, and is important for large-scale applications such as batch scanning of software. According to the detection mechanism, Perplexity applies perturbation strategies to generate dozens of perturbation samples and uses an LLM to calculate the probability or rank of each token for both the original sample and perturbation samples (Shi et al., 2025; Su et al., 2023). Such a two-phase detection process might be time-consuming. In $\text{RQ}_{2}$, we compare Perplexity with feature- and pre-training-based methods regarding detection speed.
  • $\text{{RQ}}_{\textbf{3}}$:
@@
    Developers utilize various LLMs to assist in programming. Due to differences in training data, model architecture, and generation strategies, the code generated by different LLMs varies significantly. An optimal detection method should be capable of detecting the code generated by different LLMs. In $\text{RQ}_{3}$, we investigate the generalization capability of Perplexity in detecting LLMgCode.
4.2. Data Preparation
Although there are datasets (Shi et al., 2025; Bukhari et al., 2023; Idialu et al., 2024) that work to detect LLMgCode from HaCode, none of them is suitable to address all research questions. These datasets have the following limitations: (1) they support only Python and C (limited diversity in languages); (2) their LLMgCode code is generated from a single LLM or a single family of LLMs (limited capability in generalization); and (3) only one dataset (Idialu et al., 2024) provides difficulty scores of puzzles, (limited degrees in difficulty). Therefore, we developed a new dataset based on CodeNet (Puri et al., 2021). CodeNet was developed by IBM to promote AI technologies in the code domain and has been widely used for many studies (Xu and Sheng, 2024; Agarwal et al., 2021; Goel et al., 2022; Pan et al., 2024; Naik et al., 2022, 2023). CodeNet comprises a large number of submissions from real-world developers. More than half of these submissions are functionally correct (accepted), and more than 90% are successfully compiled and executed (executable). The high quality of CodeNet ensures our HaCode serves as a reliable baseline.
Figure 5. An example of a programming puzzle in CodeNet
@@ 4.2. Data Preparation
As shown in Figure 5, a programming puzzle provided by CodeNet has seven basic elements: (1) difficulty score, (2) statement, (3) constraint, (4) input, (5) output, (6) sample input, and (7) sample output. The difficulty score measures the difficulty level of the puzzle. The statement describes what the puzzle is. The constraint describes the specific constraints on the variables. The input and output describe the requirements for input and output, respectively. Finally, the sample input and sample output provide the input and output of a specific example, respectively. Based on CodeNet, we developed our dataset through the following five steps.
@@
Figure 6. The prompts used for generating LLMgCode solutions for programming puzzles
  • Step-1: Problem selection. To build an appropriate dataset for evaluation, we excluded the programming problem whose (1) statement is not in English as it is the most popular language; (2) statement contains images or tables as they can hardly be extracted and input into LLMs; (3) solutions are all less than 30 LOC (Lines Of Code) as they are mainly for trivial puzzles; (4) solutions are incomplete regarding all the eight programming languages as they constitute the HaCode to be evaluated in $\text{RQ}_{1.1}$; and (5) structure is not complete, i.e., statement, input, output, constraints, difficulty, and samples ($\geq 2$). After employing these criteria, our dataset includes 729 programming puzzles.
  • Step-2: HaCode collection. For each puzzle selected in Step-1, we randomly took two code solutions for each of the eight programming languages from CodeNet. The code solutions were submitted by human developers, including correct, incorrect, and rejected solutions. Finally, we had 16 HaCode solutions for each puzzle.
  • Step-3: LLMgCode generation. We generated LLMgCode for programming puzzles using OpenAI GPT-3.5 (gpt-3.5-turbo-0125). The prompts for code generation are shown in Figure 6. There are two solutions for each puzzle regarding each of the eight programming languages. Moreover, to address $\text{RQ}_{3}$, we used the prompts from Figure 6 and input them into GPT-4o, Gemini-1.0, and Llama-3.1, with each model generating 500 solutions. Note that the following two steps are performed on the code solutions that have been generated by GPT-3.5.
  • Step-4: Classification according to the difficulty of puzzles. We categorized the puzzles based on their difficulty degrees and set two thresholds for the difficulty score to ensure that each group has the same number of puzzles. The detailed settings are: easy for $\leq 200$, medium for $>200$ and $\leq 500$, and hard for $>500$.
  • Step-5: Classification according to the scale of solutions. We categorized the code solutions based on their scale (LOC) and set two thresholds to ensure that each group has the same number of solutions. The detailed settings are: small for $\leq 20$, medium for $>20$ and $\leq 50$, and large for $>50$.
Finally, we obtained a dataset containing function-level LLMgCode and HaCode that solve the same programming puzzles, which is used for the experimental evaluations. We perform the evaluation on function-level code because (1) due to the current limitations of LLMgCode in handling complex programming tasks, developers usually use it to assist in implementing functions rather than entire software repositories. As a result, repo-level detection also needs to be performed at the function level, and (2) to the best of our knowledge, existing LLMgCode detection methods are all designed for function-level detection. Therefore, we evaluate them using the function-level code to ensure a fair evaluation.
4.3. Studied Methods
We compared Perplexity against feature- and pre-training-based methods.
4.3.1. Feature-based Methods
Feature-based methods require professionals to devise appropriate features of code, e.g., lexical and syntactic features, and input these features into machine learning classifiers for detection. We investigate two feature-based methods as follows.
  • Idialu et al. (Idialu et al., 2024) takes a combination of layout, syntactic, and lexical features of Python programs as the input, where layout features refer to code formatting such as indentation and spacing, lexical features are derived from code tokens such as keywords and literals, and syntactic features capture the structural patterns and relationships between code elements. Idialu uses XGBoost to construct the detection method.